OpenStudy (anonymous):

Square root of 45x^3y^3

OpenStudy (saifoo.khan):

\[\Huge \sqrt{45}x^\frac{3}{2}y^\frac{3}{2}\]

jimthompson5910 (jim_thompson5910):

\[\sqrt{45x^3y^3}\] \[\sqrt{9*5x^2*x*y^2*y}\] \[\sqrt{9}*\sqrt{5}*\sqrt{x^2}*\sqrt{x}*\sqrt{y^2}*\sqrt{y}\] \[3*\sqrt{5}*x*\sqrt{x}*y*\sqrt{y}\] \[3xy\sqrt{5xy}\] So \[\sqrt{45x^3y^3}=3xy\sqrt{5xy}\]

jimthompson5910 (jim_thompson5910):

you can go further with simplifying saifoo